# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v1

@@ Abstract
In recent years, Large-Language-Model-driven AI agents have exhibited unprecedented intelligence, flexibility, and adaptability, and are rapidly changing human production and lifestyle. Nowadays, agents are undergoing a new round of evolution. They no longer act as an isolated island like LLMs. Instead, they start to communicate with diverse external entities, such as other agents and tools, to collectively perform more complex tasks. Under this trend, agent communication is regarded as a foundational pillar of the future AI ecosystem, and many organizations intensively begin to design related communication protocols (e.g., Anthropic’s MCP and Google’s A2A) within the recent few months. However, this new field exposes significant security hazard, which can cause severe damage to real-world scenarios. To help researchers to quickly figure out this promising topic and benefit the future agent communication development, this paper presents a comprehensive survey of agent communication security. More precisely, we first present a clear definition of agent communication and categorize the entire lifecyle of agent communication into three stages: user-agent interaction, agent-agent communication, and agent-environment communication. Next, for each communication phase, we dissect related protocols and analyze its security risks according to the communication characteristics. Then, we summarize and outlook on the possible defense countermeasures for each risk. Finally, we discuss open issues and future directions in this promising research field.
@@
large language model, AI agent, agent communication, attack and security
@@ I Introduction
The emergence of Large Language Models (LLMs) has led to revolutionary advancements in Artificial Intelligence (AI), exhibiting unprecedented capabilities in understanding complex tasks [308]. More importantly, LLMs greatly boosted the ideal form of AI that human expects: agents [^1] In this paper, all agents refer to LLM-driven AI agents.. Different from LLMs that are mainly like chat bots, agents possess more comprehensive capabilities (e.g., perception, interaction, reasoning, and execution), enabling them to independently complete a real-world task. For example, when users seek to make a travel plan, LLMs can only provide the best plan in text, while agents can realize the best plan in action, such as checking the weather, buying tickets, and booking hotels. Agents greatly speed up the progress of the intelligence transformations of enterprise. Their market size is respected to increase by 46% per year [222]. It can be foreseen that agents will subvert the production and living patterns of modern society, greatly changing the future business landscape. As a result, developing and promoting agents have become the strategic planning of major powers and influential companies.
@@
Figure 1: The comparison between traditional Internet and Internet of Agents (IoA). In traditional Internet, users need to manually visit different websites to finish a travel, which is cumbersome. With Internet of Agents, users only need to assign a task to their agent, which will communicate with the agents of different companies (e.g., hotel and train companies) to automatically finish a best travel plan.
@@
Now, agents are evolving towards the direction of domain-specific entities, i.e., being customized for specific scenarios and tasks. In this context, as shown in Figure 1, a task usually requires the collaboration of multiple agents, which may locate globally on the Internet. Under this condition, agent communication becomes the foundation of the future AI ecosystem. It enable agents to find other agents with specific capabilities, access external knowledge, assign tasks, and achieve other interactions. Based on the vast market of agent communication, an increasing number of communities and companies are seizing the opportunity to contribute to the development of agent communication. In November, 2024, Anthropic proposed Model Context Protocol (MCP) [16], a universal protocol that allows agents to invoke external environments, such as datasets, tools, and APIs. MCP quickly gained a great deal of attention in the recent few months. Up to now, hundreds of enterprises have announced their access to MCP, such as OpenAI [203], Google [87], Microsoft [53], Amazon [21], Alibaba [10], and Tencent [251], and MCP’s package receives over 3 million weekly downloads [17]. In April 2025, Google proposed Agent to Agent Protocol (A2A) [218], which enables seamless communication and collaboration among agents. Since its release, A2A has received extensive supports from many enterprises, such as Microsoft [188], Atlassian [149], and PayPal [229]. It can be seen that the breakthroughs in agent communication are bringing rapid and profound changes and will become an indispensable part of the AI ecosystem.
@@
However, the rapid development of agent communication also introduce complex security risks that could cause severe damage in the AI ecosystem. For example, the collaboration of cross-organization agents significantly enlarging the attack surface, causing severe security risks, including but not limited to privacy leakage, agent spoofing, agent bullying, and Denial of Service attacks. Since the research related to agent communication is still in the nascent stage, it urgently needs a systematic review of the security problems existing in the complete agent communication lifecycle. Following this trend, this paper aims to provide a comprehensive survey of existing agent communication techniques, analyze their security risks, and discuss possible defense countermeasures. We believe this work can help a broad range of readers, such as researchers who devote to agent development and beginners who just start their journey in AI.
@@
  • We present the first systematic overview of agent communication. Specifically, we propose a definition of agent communication for the first time, and classify it into three stages based on the communication object: user-agent interaction, agent-agent communication, and agent-environment communication. This classification covers the entire lifecycle of agent communication, and the protocols in the same stage exhibit similar attack surfaces, which can help future studies to analyze and evaluate their work more conveniently.
  • We make in-depth analyses of the security risks in the evolvement of agent communication, discussing the discovered attacks and those that have yet not been revealed. Our analyses show that user-agent interaction mainly suffers from threats from faulty and malicious user inputs, agent-agent communication faces significant attacks from other agents and in-the-middle adversaries, and agent-environment communication can be easily impacted by compromised external tools and resources.
  • We detailedly discuss the targeted defense countermeasures for mitigating the exposed security risks, pointing out the possible development directions. For user-agent interaction, effective filter of multimodal inputs is necessary; agent-agent communication needs powerful mechanisms to monitor, archive, audit, and quantify the responsibility of actions in agent collaboration; agent-environment communication should be protected by powerful detection of poisoned external environments.
  • We finally discuss the open issues and future research directions. We not only point out the much needed techniques that can defend agent communication, but also explained that related laws and regulations should be improved as soon as possible. Only when both aspects develop simultaneously can the security of agent communication in practice gets guaranteed.
@@
Organization. As shown in Figure 2, we organize this survey as follows. Section II compares the most relevant surveys with this paper and outlines the novelties in this survey. Section III introduces the preliminaries of this survey. Section IV presents a definition and classification of agent communication. Section V introduces user-agent interaction protocols and analyzes related security risks and defense countermeasures. Section VI exhibits agent-agent communication protocols, related security risks, and corresponding defense countermeasures. Similarly, Section VII shows the protocols, risks, and defenses for user-environment communication. In Section VIII, we discuss the open issues and future research direction. Section IX concludes this survey.
@@ II Related Work
TABLE I: Comparison between different surveys, where “Moti.” refers to the Motivation of proposing agent communication; “Defi.” refers to the Definition of agent communication; “Clas.” refers to the Classification of agent communication or protocols; “U-A” refers to user-agent Interaction; “A-A” refers to agent-agent communication; “A-E” refers to agent-environment communication; “Secu.” refers to Security; “Comm.” refers to Communication; “Research Object” denotes the theme of a survey; “Agent Communication” denotes whether a survey concentrate on agent communication; “Protocol Coverage” denotes whether a survey includes comprehensive agent communication protocols; “Security Analyses” denotes whether a survey analyzes the security risks of different agent communication stages; “Defense Prospect” denotes whether a survey analyzes the possible defenses for different agent communication stage; “Rele.” refers to the degree of relevance between a survey and this survey, where the higher the score, the more relevant it is; : Not discussed in this survey; : Mentioned But not a main focus or not discussed comprehensively in this survey; : comprehensively discussed in this survey.
Survey Year Research Object Rele. Agent Communication Protocol Coverage Security Analyses Defense Prospect
@@ II-A Overview of Novelties in This Survey
  • This survey presents a comprehensive illustration of agent communication. Specifically, it explains why the current agent ecosystem needs communication (i.e., the predicament faced by a single agent, Section IV-A), gives the definition of agent communication (Section IV-B), and proposes a novel classification principle based on communication entity, which can cover the entire lifecyle of agent communication (Section IV-C). As a result, future studies can be included and categorized according to our survey seamlessly, which benefits the systematic study and development of this field.
  • This survey exhibits a comprehensive illustration of the existing protocols related to different agent communication stages (Sections V-A, VI-A, and VII-A), including newly proposed and previously neglected protocols that have not been discussed by other surveys. Besides, we categorize these protocols based on their architecture and summarize corresponding characteristics, rather than mechanically listing each protocol. This organization method can allow any researchers interested in this field quickly establish a preliminary but comprehensive understanding of agent communication.
  • This survey makes an in-depth analysis of the found and potential security risks that have not been revealed for each agent communication stage (Sections V-B, VI-B, and VII-B). We introduce both the vulnerability from which attacks derive and typical attack examples, systematically pointing out the drawbacks of the current agent communication. To our knowledge, there have not been studies mainly focusing on the security risks of agent communication.
  • This survey throughly outlooks on the possible defense countermeasures (Sections V-C, VI-C, and VII-C) that can make future agent communication more secure and reliable, covering the whole lifecycle of agent communication, which is not achieved by previous surveys.
@@ II-B Selection Principles of the Most Relevant Surveys
To solve this challenge, when selecting the most relevant surveys, we focus on three principles: LLM-driven agents, agent communication, and security. However, to our knowledge, there have not been papers systematically discuss all of these three themes. As a result, as long as a survey meets two of the three principles, we will treat it as a relevant survey.
  • Principle #1: LLM-driven agents. The first and the most important is that the research object of a survey must be LLM-driven agents. This principle must be satisfied. This is because there have been many studies about multi-agent systems (MAS) before the emergence of LLMs. These agents have completely different cores and characteristics from LLM-driven agents, so discussing them benefits very little to this survey. Besides, surveys focusing on only LLMs instead of LLM-driven agents are also not listed in Table I (but we will draw on their valuable insights in other sections). This is because LLMs shows significant difference from agents, which we make a detailed illustration in Section III-C. As a result, researching LLM-driven agents is the most important principle.
  • Principle #2: agent communication. The second principle is that a survey focus on or partially discuss agent communication, especially including some typical agent communication protocols such as MCP. This is because agent communication is very different from agent. However, if a survey satisfies the other two principles (i.e., LLM-driven agents and security), we still treat it as a relevant survey.
  • Principle #3: security. The final principle is that a survey focus or partially discussed agent-related security. This is because we believe that the security risks of agents still have meaning to the security risks of agent communication. The former is usually a subset of the latter, i.e., agent communication shows novel and more attack surfaces compared to agent.
Relevance Score. As a result, we can find that there are two main types of relevant surveys: LLM-driven agents + communication, or LLM-driven agents + security. As shown in Table I, we list a relevance score for each survey. The higher the score, the more relevant we think it is to our survey. This score is subjectively derived by us after carefully reading the paper and does not have an objective calculation method. This is because we found that the forms of surveys are highly diverse, and it is hard to accurately classify them using merely several metrics. As a result, we directly present the score based on our subjective feelings when reading these surveys.
@@ II-C Detailed Comparison with the Most Relevant Surveys
Survey [161] focuses on personal agents that deeply integrate personal data and devices, exploring their potential as the main software paradigm for future personal computer. It only partially mentions the security risks related to personal agents in a section. Besides, these risks only belongs to the user-agent interaction phase. It also does not discuss agent communication and related security risks and defenses. Survey [72] focuses on agent security instead of agent communication security. It is a security-specific paper, so its discussion of security is more comprehensive compared to [161]. However, the main body of its discussion is about the interaction between user and agent (U-A), without enough considerations about agent-agent (A-A) or agent-environment (A-E) interaction, which have significantly different characteristics. Besides, it also does not include any protocols related to agent communication. Survey [153] also focuses on agent security instead of agent communication security, which is similar to [72]. This survey focuses on single-agent systems and partially discussed multi-agent collaboration. It does not consider agent communication, related protocols, and enough security analyses about A-A and A-E. Survey [133] systematically summarizes seven security challenges for multi-agent systems. As shown in Table I, its main focus in on A-A, and only partially discusses U-A and A-E, which is not comprehensive. Besides, it does not consider agent communication and related protocols. Survey [52] propose four knowledge gaps faced by agents, which mainly fall within U-A, partially discussing A-A and A-E. Besides, it does not consider agent communication or any related protocols. The defense prospect is also limited. Survey [98] focuses on the security risks of U-A and A-E, such as malicious API. It does not consider agent communication and related protocols. Besides, its security analyses are also not comprehensive enough. Survey [276] focus on the fundamentals, applications, and challenges of IoA. Since its focus is different, agent communication and related security are only partially mentioned. Specifically, it only introduces a few related protocols and briefly makes an analysis about related security. Besides, it also lacks the related illustration (such as definition and classification) of agent communication. Survey [93] also concentrates on the security of U-A, partially discussing A-E. It does not mention agent communication and related protocols, as well as the risks of A-A. Survey [298] focus on agent communication architecture, which is a study of agent interaction mechanisms from a high-level and abstract view. Besides, it only partially mentioned related security, and did not discuss any communication protocols. Survey [277] focuses on the security of IoA. It mentioned a few agent communication protocols (i.e., MCP, A2A, ANP, and Agora), neglecting many other important protocols. Besides, it lacks the motivation, definition, and classification of agent communication, and also does not classify protocols. According to our analyses, the security analyses (especially for U-A) are also not comprehensive enough. Survey [263] propose the concept of “full stack safety” of agents, providing comprehensive analyses of data preparation, pre-training, post-training, deployment, and commercialization. It does not focus on agent communication security. As a result, this survey did not give a clear illustration of agent communication, only mentioned a few protocols (i.e., MCP, A2A, ANP, and Agora), and partially discussed related threats and countermeasures. Survey [278] gives a comprehensive analyses of the security of agent networks. However, it does not include the discussion of communication protocols, and lacks enough security analyses of A-A and A-E. Survey [69] does not focus on security. Instead, it concentrates on evaluating LLMs and agents. Besides, it also analyzes the architecture of some communication protocols (i.e., MCP, A2A, and ANP). We can see that it does not give a detailed illustration of agent communication, enough coverage of protocols, or comprehensive discussion about security. Survey [228] focuses on MCP, detailedly analyzed related architectures and applications. It does not consider other communication protocols and only partially mentioned security-related contents. Survey [306] analyzed the threats of agents and dividing them into two categories (intrinsic and extrinsic), partially covering U-A, A-A, and A-E. However, its analyses are not comprehensive enough, and it did not mention any communication protocols. Survey [86] makes a comprehensive analyses of the risks for multi-agent systems. However, its focus only fall within A-A, not considering U-A, A-E, and related protocols.
@@
Survey [299] is one of the surveys with highest relevance scores because it focuses on agent communication and analyzes related protocols. However it still have significant differences from our survey. First, it lacks some critical protocols like AG-UI, ACP-AgentUnion, ACN, Agent Protocol, API Bridge Agent, and Function Calling. Second, it lacks analyses of security threats. Third, its defenses prospect is limited. Survey [137] focuses on the influences of MCP. As a result, it lacks other protocols, the illustration of agent communication, and security-related discussion. Similarly, survey [239] also focuses on MCP. It lacks illustrations of other protocols and comprehensive security analyses. Survey [220] comprehensively introduces A2A, lacking discussion about other protocols and security analyses. Survey [62] detailed discussed MCP, A2A, ANP, and ACP(-IBM). It also partially analyzed related security risks and defenses. However, there still lacks other protocols, in-depth security analyses, and systematic illustration of agent communication. Hou et al. [101] discussed the security risks of MCP. They did not consider other protocols and the high-level overview of agent communication.

### v2

@@ Abstract
In recent years, Large-Language-Model-driven AI agents have exhibited unprecedented intelligence and adaptability, and are rapidly changing human production and life. Nowadays, agents are undergoing a new round of evolution. They no longer act as an isolated island like LLMs. Instead, they start to communicate with diverse external entities, such as other agents and tools, to perform more complex tasks collectively. Under this trend, agent communication is regarded as a foundational pillar of the future AI ecosystem, and many organizations have intensively begun to design related communication protocols (e.g., Anthropic’s MCP and Google’s A2A) within the recent few months. However, this new field exposes significant security hazards, which can cause severe damage to real-world scenarios. To help researchers quickly figure out this promising topic and benefit the future agent communication development, this paper presents a comprehensive survey of agent communication security. More precisely, we first present a clear definition of agent communication and categorize the entire lifecycle of agent communication into three stages: user-agent interaction, agent-agent communication, and agent-environment communication. Next, for each communication phase, we dissect related protocols and analyze the security risks according to the communication characteristics. Then, we summarize and outlook on the possible defense countermeasures for each risk. In addition, we conduct experiments using MCP and A2A to help readers better understand the novel vulnerabilities brought by agent communication. Finally, we discuss open issues and future directions in this promising research field.
@@
large language model, agent communication, attack, and security
@@ I Introduction
The emergence of Large Language Models (LLMs) has led to revolutionary advancements in Artificial Intelligence (AI), exhibiting unprecedented capabilities in understanding complex tasks [327]. More importantly, LLMs greatly boosted the ideal form of AI that human expects: agents [^1] In this paper, all agents refer to LLM-driven AI agents.. Different from LLMs that are mainly like chatbots, agents possess more comprehensive capabilities (e.g., perception, interaction, reasoning, and execution), enabling them to independently complete a real-world task. For example, when users seek to make a travel plan, LLMs can only provide the best plan in text, while agents can realize the best plan in action, such as checking the weather, buying tickets, and booking hotels. Agents greatly speed up the progress of the intelligence transformations of enterprises. Their market size is expected to increase by 46% per year [235]. It can be foreseen that agents will subvert the production and living patterns of modern society, greatly changing the future business landscape. As a result, developing and promoting agents have become a strategic planning of major powers and influential companies.
@@
Figure 1: The comparison between traditional Internet and Internet of Agents (IoA). In the traditional Internet, users need to manually visit different websites to finish a trip, which is cumbersome. With Internet of Agents, users only need to assign a task to their agent, which will communicate with the agents of different companies (e.g., hotel and train companies) to automatically finish the best travel plan.
@@
Now, agents are evolving towards the direction of domain-specific entities, i.e., being customized for specific scenarios and tasks. In this context, as shown in Figure 1, a task usually requires the collaboration of multiple agents, which may be located globally on the Internet. Under this condition, agent communication becomes the foundation of the future AI ecosystem. It enables agents to find other agents with specific capabilities, access external knowledge, assign tasks, and engage in other interactions. Based on the vast market of agent communication, an increasing number of communities and companies are seizing the opportunity to contribute to the development of agent communication. In November 2024, Anthropic proposed Model Context Protocol (MCP) [19], a universal protocol that allows agents to invoke external environments, such as datasets, tools, and APIs. MCP quickly gained a great deal of attention in the recent few months. Up to now, hundreds of enterprises have announced their access to MCP, such as OpenAI [214], Google [96], Microsoft [57], Amazon [25], Alibaba [10], and Tencent [267], and MCP’s package receives over 3 million weekly downloads [21]. In April 2025, Google proposed Agent-to-Agent Protocol (A2A) [89], which enables seamless communication and collaboration among agents. Since its release, A2A has received extensive support from many enterprises, such as Microsoft [200], Atlassian [160], and PayPal [242]. It can be seen that the breakthroughs in agent communication are bringing rapid and profound changes and will become an indispensable part of the AI ecosystem.
@@
However, the rapid development of agent communication also introduces complex security risks that could cause severe damage to the AI ecosystem. For example, the collaboration of cross-organization agents significantly enlarges the attack surface, causing severe security risks, including but not limited to privacy leakage, agent spoofing, agent bullying, and Denial of Service attacks. Since the research related to agent communication is still in the nascent stage, it urgently needs a systematic review of the security problems existing in the complete agent communication lifecycle. Following this trend, this paper aims to provide a comprehensive survey of existing agent communication techniques, analyze their security risks, and discuss possible defense countermeasures. We believe this work can help a broad range of readers, such as researchers who are devoted to agent development and beginners who have just started their journey in AI.
@@
  • We propose the definition of agent communication for the first time, and classify it into three stages based on the communication object, which can cover the entire lifecycle of agent communication.
  • We make in-depth analyses of the security risks in agent communication, discussing the discovered attacks and those that have yet not be revealed.
  • We detailedly discuss the targeted defense countermeasures for mitigating the exposed security risks.
  • We conduct experiments using MCP and A2A, clearly illustrating the novel attack surfaces brought by agent communication.
  • We finally discuss the open issues and future research directions. We not only point out the much-needed techniques but also explain the demand for related laws and regulations.
@@
Organization. As shown in Figure 2, we organize this survey as follows. Section II compares the most relevant surveys with this paper and outlines the novelties in this survey. Section III introduces the preliminaries of this survey. Section IV presents a definition and classification of agent communication. Section V introduces user-agent interaction protocols and analyzes related security risks and defense countermeasures. Section VI exhibits agent-agent communication protocols, related security risks, and corresponding defense countermeasures. Similarly, Section VII shows the protocols, risks, and defenses for agent-environment communication. In Section VIII, we conduct experiments using MCP and A2A to help illustrate the risks brought by agent communication. In Section IX, we discuss the open issues and future research direction. Section X concludes this survey.
@@ II Related Work
TABLE I: Comparison between different surveys, where “Moti.” refers to the Motivation of proposing agent communication; “Defi.” refers to the Definition of agent communication; “Clas.” refers to the Classification of agent communication or protocols; “U-A” refers to user-agent Interaction; “A-A” refers to agent-agent communication; “A-E” refers to agent-environment communication; “Secu.” refers to Security; “Comm.” refers to Communication; “Research Object” denotes the theme of a survey; “Agent Communication” denotes whether a survey concentrate on agent communication; “Protocol Coverage” denotes whether a survey includes comprehensive agent communication protocols; “Security Analyses” denotes whether a survey analyzes the security risks of different agent communication stages; “Defense Prospect” denotes whether a survey analyzes the possible defenses for different agent communication stage; “Exp.” refers to the experiments of attacks in agent communication; “Rele.” refers to the degree of relevance between a survey and this survey, where the higher the score, the more relevant it is; : Not discussed in this survey; : Mentioned But not a main focus or not discussed comprehensively in this survey; : comprehensively discussed in this survey.
Survey Year Research Object Rele. Agent Communication Protocol Coverage Security Analyses Defense Prospect Exp.
@@ II-A Overview of Novelties in This Survey
  • This survey presents a comprehensive illustration of agent communication. Specifically, it explains why the current agent ecosystem needs communication (i.e., the predicament faced by a single agent, Section IV-A), gives the definition of agent communication (Section IV-B), and proposes a novel classification principle based on communication entity, which can cover the entire lifecyle of agent communication (Section IV-C). As a result, future studies can be included and categorized according to our survey seamlessly.
  • This survey exhibits a comprehensive illustration of the existing protocols related to different agent communication stages (Sections V-A, VI-A, and VII-A), including newly proposed and previously neglected protocols that have not been discussed by other surveys. Besides, we categorize these protocols based on their architecture and summarize corresponding characteristics, rather than mechanically listing each protocol. This organization method can allow any researchers interested in this field to quickly establish a preliminary but comprehensive understanding of agent communication.
  • This survey makes an in-depth analysis of the discovered attacks and potential risks that have not been revealed for each agent communication stage. We discuss both the risks from compromised agents and those from malicious communication objects (i.e., users and environments), fully covering the entire agent communication lifecycle. Then, we thoroughly outline the possible defense countermeasures (Sections V-D, VI-C, and VII-D) that can make future agent communication more secure.
  • This survey conducts experiments using MCP and A2A (Section VIII), the most popular agent communication protocols up to now. We successfully launched attacks against MCP and A2A, showing that attackers can cause severe damage with little effort. This section can help readers better understand the new attack surfaces brought by agent communication.
@@ II-B Selection Principles of the Most Relevant Surveys
To solve this challenge, when selecting the most relevant surveys, we focus on three principles: LLM-driven agents, agent communication, and security. However, to our knowledge, there have not been papers systematically discussing all three of these themes. As a result, as long as a survey meets two of the three principles, we will treat it as a relevant survey.
  • Principle #1: LLM-driven agents. The first and the most important is that the research object of a survey must be LLM-driven agents. This principle must be satisfied. This is because there have been many studies about multi-agent systems (MAS) before the emergence of LLMs. These agents have completely different cores and characteristics from LLM-driven agents, so discussing them benefits very little to this survey. Besides, surveys focusing on only LLMs instead of LLM-driven agents are also not listed in Table I (but we will draw on their valuable insights in other sections). This is because LLMs show significant differences from agents, for which we make a detailed illustration in Section III-C. As a result, researching LLM-driven agents is the most important principle.
  • Principle #2: agent communication. The second principle is that a survey should focus on or partially discuss agent communication, especially including some typical agent communication protocols such as MCP. This is because agent communication is very different from agent. However, if a survey satisfies the other two principles (i.e., LLM-driven agents and security), we still treat it as a relevant survey.
  • Principle #3: security. The final principle is that a survey should focus on or partially discuss agent-related security. This is because we believe that the security risks of agents still have meaning to the security risks of agent communication. The former is usually a subset of the latter, i.e., agent communication shows novel and more attack surfaces compared to agent.
Relevance Score. As a result, we can find that there are two main types of relevant surveys: LLM-driven agents + communication, or LLM-driven agents + security. As shown in Table I, we list a relevance score for each survey. The higher the score, the more relevant we think it is to our survey. This score is subjectively derived by us after carefully reading the paper and does not have an objective calculation method. This is because we found that the forms of surveys are highly diverse, and it is hard to accurately classify them using only several metrics. As a result, we directly present the score based on our subjective feelings when reading these surveys.
@@ II-C Detailed Comparison with the Most Relevant Surveys
Survey [170] focuses on personal agents that deeply integrate personal data and devices, exploring their potential as the main software paradigm for future personal computers. It only partially mentions the security risks related to personal agents in a section. Besides, these risks only belong to the user-agent interaction phase. It also does not discuss agent communication and related security risks and defenses. Survey [77] focuses on agent security instead of agent communication security. It is a security-specific paper, so its discussion of security is more comprehensive compared to [170]. However, the main body of its discussion is about the interaction between user and agent (U-A), without enough consideration about agent-agent (A-A) or agent-environment (A-E) interaction, which have significantly different characteristics. Besides, it also does not include any protocols related to agent communication. Survey [163] also focuses on agent security instead of agent communication security, which is similar to [77]. This survey focuses on single-agent systems and partially discusses multi-agent collaboration. It does not consider agent communication, related protocols, and enough security analyses about A-A and A-E. Survey [143] systematically summarizes seven security challenges for multi-agent systems. As shown in Table I, its main focus is on A-A, and only partially discusses U-A and A-E, which is not comprehensive. Besides, it does not consider agent communication and related protocols. Survey [56] proposes four knowledge gaps faced by agents, which mainly fall within U-A, partially discussing A-A and A-E. Besides, it does not consider agent communication or any related protocols. The defense prospect is also limited. Survey [108] focuses on the security risks of U-A and A-E, such as malicious API. It does not consider agent communication and related protocols. Besides, its security analyses are also not comprehensive enough. Survey [295] focuses on the fundamentals, applications, and challenges of IoA. Since its focus is different, agent communication and related security are only partially mentioned. Specifically, it only introduces a few related protocols and briefly analyzes related security. Besides, it also lacks the related illustration (such as definition and classification) of agent communication. Survey [103] also concentrates on the security of U-A, partially discussing A-E. It does not mention agent communication and related protocols, as well as the risks of A-A. Survey [317] focuses on agent communication architecture, which is a study of agent interaction mechanisms from a high-level and abstract view. Besides, it only partially mentioned related security and did not discuss any communication protocols. Survey [296] focuses on the security of IoA. It mentioned a few agent communication protocols (i.e., MCP, A2A, ANP, and Agora), neglecting many other important protocols. Besides, it lacks the motivation, definition, and classification of agent communication, and also does not classify protocols. According to our analyses, the security analyses (especially for U-A) are also not comprehensive enough. Survey [281] proposes the concept of “full stack safety” of agents, providing comprehensive analyses of data preparation, pre-training, post-training, deployment, and commercialization. It does not focus on agent communication security. As a result, this survey did not give a clear illustration of agent communication, only mentioned a few protocols (i.e., MCP, A2A, ANP, and Agora), and partially discussed related threats and countermeasures. Survey [297] gives a comprehensive analysis of the security of agent networks. However, it does not include the discussion of communication protocols, and lacks enough security analyses of A-A and A-E. Survey [75] does not focus on security. Instead, it concentrates on evaluating LLMs and agents. Besides, it also analyzes the architecture of some communication protocols (i.e., MCP, A2A, and ANP). We can see that it does not give a detailed illustration of agent communication, enough coverage of protocols, or a comprehensive discussion about security. Survey [241] focuses on MCP, detailedly analyzed related architectures and applications. It does not consider other communication protocols and only partially mentions security-related content. Survey [325] analyzes the threats of agents and divides them into two categories (intrinsic and extrinsic), partially covering U-A, A-A, and A-E. However, its analyses are not comprehensive enough, and it does not mention any communication protocols. Survey [95] makes a comprehensive analysis of the risks for multi-agent systems. However, its focus only falls within A-A, not considering U-A, A-E, and related protocols.
@@
Survey [318] is one of the surveys with the highest relevance score because it focuses on agent communication and analyzes related protocols. However, it still has significant differences from our survey. First, it lacks some critical protocols like AG-UI, ACP-AgentUnion, ACN, Agent Protocol, API Bridge Agent, and Function Calling. Second, it lacks the analysis of security threats. Third, its defenses prospect is limited. Survey [147] focuses on the influences of MCP. As a result, it lacks other protocols, the illustration of agent communication, and security-related discussion. Similarly, the survey [252] also focuses on MCP. It lacks illustrations of other protocols and comprehensive security analyses. Survey [233] comprehensively introduces A2A, lacking discussion about other protocols and security analyses. Survey [67] detailed discussed MCP, A2A, ANP, and ACP(-IBM). It also partially analyzed related security risks and defenses. However, there still lacks other protocols, in-depth security analyses, and systematic illustration of agent communication. Hou et al. [111] discussed the security risks of MCP. They did not consider other protocols and the high-level overview of agent communication.